# Versatile Backdoor Attack with Visible, Semantic, Sample-specific and Compatible Triggers

### v3

@@ V-C1 Experimental settings
In the object detection task, we introduce two different attack types following BadDet [22], Object Disappearance Attack (ODA) and Global Misclassification Attack (GMA). ODA aims to make all bounding boxes in poisoned images vanish, while GMA aims to misclassify all bounding boxes in poisoned images as the target class. We only introduce fundamental settings in the following sections. For further details, please refer to Appendix A-B.
@@
Dataset and models. The dataset used in the object detection task is VOC 07+12 [65], which contains 19,352 images of 20 classes of objects. We use YOLO-v4 as the backbone model, which is a widely used anchor-based one-stage detection model. The input size of the YOLO-v4 model is 3 $\times$ 416 $\times$ 416.
@@
Implementation details. In ODA, the width and height of the poisoned bounding box are set to 0, i.e., $g_{y}([a,b,w,h,c])=[a,b,0,0,c]$. In GMA, the class of the poisoned bounding box is set to the target class, i.e., $g_{y}([a,b,w,h,c])=[a,b,w,h,y_{t}]$, we randomly select “aeroplane” as the target class with a target label $y_{t}=0$ assigned. $(a,b)$ denotes the coordinate of the top-left point of the bounding box. The poisoning ratio is set to 10% and 20%. Among all the text triggers filtered out by the trigger selection module, the ones chosen for demonstration in our experiments include “lemon”, “butterfly”, and “beachball”. For experiments in the digital-to-physical scenario, we randomly select 150 poisoned images to print and recapture. In the physical scenario, we manually capture 100 photos with the phone camera with actual trigger entities for testing.
Baseline attacks. As VSSC attack is implemented in the digital scenario, we select BadDet [22] as our baseline for comparison. In alignment with the experimental settings in BadDet, we employ a chessboard as the trigger pattern whose size and position are consistent with the original description.
@@
Evaluation metrics. For the object detection task, we still use the same evaluation metrics as the image classification task, ASR, C-Acc, and R-Acc. The definition of ASR in this section varies according to different attack types. For ODA, ASR is defined as the proportion of original bounding boxes that successfully vanish. In GMA, ASR is defined as the proportion of non-target class bounding boxes that are misclassified as the target class. C-Acc is defined as mAP on benign images, which represents the mean of average precision (AP) of each class. AP represents the area under the precision-recall curve for bounding boxes with a sufficient confidence level. Our experiments use mAP on an IoU threshold of 0.5 (mAP@0.5). R-Acc is defined as mAP on poisoned images. ASR and R-Acc do not reach a high level simultaneously, but there is no additive relationship between them.
Figure 5: Poisoned samples created by BadNets and VSSC for the object detection task. For the VSSC attack, we showcase three triggers: “lemon”, “butterfly”, and “beachball”. We only use the results under Global Misclassification Attack (GMA) for demonstration, where D2P means the digital-to-physical scenario.
@@ V-C2 Effectiveness in digital scenario
As shown in Table V, in the digital scenario, for ODA task, our attack achieves an ASR up to 99.03% with a poisoning ratio of 20% and an ASR up to 97.61% with a poisoning ratio of 10%. For GMA task, our attack achieves an ASR up to 86.49% with a poisoning ratio of 20%, and an ASR up to 84.76% with a poisoning ratio of 10%. Despite using sample-specific triggers, the VSSC attack outperforms BadDet under different attack types and poisoning ratios, achieving a higher ASR and a smaller C-Acc decrease. This demonstrates the effectiveness of the VSSC attack in the digital scenario for object detection tasks.
@@
TABLE V: Attack performance of VSSC and baseline attack in the digital scenario for the object detection task.
Attack $\downarrow$ Trigger $\downarrow$ Attack type $\rightarrow$ ODA GMA
Pratio $\downarrow$ C-Acc(%) ASR(%) R-Acc(%) C-Acc(%) ASR(%) R-Acc(%)
BadDet [22] Chessboard 10% 85.51 94.68 1.38 85.43 76.17 2.46
Chessboard 20% 85.14 97.04 1.10 84.25 84.52 0.68
VSSC (Ours) Lemon 10% 86.32 96.45 2.41 86.76 81.46 0.57
Lemon 20% 85.83 98.17 1.01 85.98 85.07 0.03
Butterfly 10% 86.99 96.67 1.66 87.19 84.06 0.64
Butterfly 20% 86.26 98.28 1.23 86.09 86.49 0.12
Beachball 10% 86.70 97.61 0.85 86.92 84.76 0.28
Beachball 20% 86.10 99.03 0.45 86.18 85.99 0.16
@@ V-C3 Effectiveness in digital-to-physical scenario
For each trigger, we randomly select 150 poisoned images, which are printed and recaptured, aiming to evaluate the robustness of our attack method in the digital-to-physical scenario. As shown in Table VI, the ASR of VSSC attack under visual distortion significantly surpasses that of BadDet. With a poisoning ratio of 10%, the VSSC trigger still achieves an ASR of over 90% in the ODA task and over 70% in the GMA task. The maximum decrease in ASR of VSSC attack is only 6.67% in ODA and 12.88% in GMA. In contrast, BadDet is almost ineffective in GMA, with most bounding boxes being classified as their original class. This result demonstrates the robustness of VSSC attack under visual distortion.
TABLE VI: Attack performance of VSSC and baseline attack in the digital-to-physical scenario for the object detection task.
Attack $\downarrow$ Trigger $\downarrow$ Attack type $\rightarrow$ ODA GMA
Pratio $\downarrow$ C-Acc(%) ASR(%) R-Acc(%) C-Acc(%) ASR(%) R-Acc(%)
BadDet [22] Chessboard 10% 82.75 80.25 53.45 84.41 27.31 68.25
Chessboard 20% 80.09 84.03 45.48 80.83 35.71 54.70
VSSC (Ours) Lemon 10% 79.90 98.29 6.78 83.02 76.92 8.76
Lemon 20% 81.61 99.15 15.27 80.35 75.64 1.94
Butterfly 10% 82.26 90.00 19.52 82.56 73.33 10.09
Butterfly 20% 81.57 92.22 19.57 79.58 77.78 7.21
Beachball 10% 81.42 92.57 23.71 81.53 71.88 11.87
Beachball 20% 81.01 94.74 18.97 82.07 73.75 8.75
@@ V-C4 Effectiveness in physical scenario
To evaluate the effectiveness of VSSC attack in the physical scenario, we capture 100 photos with corresponding trigger entities for testing. Table VII demonstrates the effectiveness of VSSC attack using a real “lemon” as the trigger. At a poisoning ratio of 20%, the ASR of the ODA task can reach 97.89%, and the ASR of the GMA task can reach 76.76%. This result indicates that the VSSC trigger can successfully mislead the object detection model in the physical scenario.
TABLE VII: Attack performance of VSSC triggers in the physical scenario for the object detection task.
Attack type Trigger Pratio C-Acc(%) ASR(%) R-Acc(%)
@@
ODA Lemon 20% 83.99 97.89 26.50
GMA Lemon 10% 83.10 69.72 14.29
GMA Lemon 20% 84.65 76.76 32.95
@@ V-D Evaluations on Face Verification Task
Figure 6: Poisoned samples after face alignment for different attacks on face verification. BadNets uses a chessboard patch as the trigger pattern, while the selected triggers for VSSC are green glasses. For comparison, Blended also uses green glasses as triggers. D2P means the digital-to-physical scenario.
@@ V-D1 Experimental settings
Face verification uses embeddings extracted from human faces to verify whether two faces belong to the same person. Backdoor attacks on face verification aim to mislead the backdoor model to identify any face as being the same as the target person when the trigger is present but keep its original performance without the trigger. In the following sections, we only introduce the basic settings. For more details, refer to Appendix A-C.
@@
Implementation details. For experiments on the face verification task, the poisoning ratio is set to 1%, and the target class is fixed at $y_{t}=0$ under an all-to-one setting. From the triggers selected by the trigger selection module, we use “green glasses” as the trigger for demonstration in our experiments. All the poisoned images are aligned before being passed to the face verification model. For experiments in the digital-to-physical scenario, we randomly select 100 poisoned images for each attack to print and recapture. For the physical scenario, 50 photos with actual trigger objects are manually captured for testing.
Baseline attacks. We use two data poisoning attack methods as our baseline attacks, BadNets [7] and Blended [8]. For BadNets, we adopt the chessboard pattern as the trigger, in line with the settings outlined in [16]. Regarding Blended, to facilitate a comparison with the VSSC attack, we also utilize green glasses as the trigger.
TABLE VIII: Attack performance of VSSC and two baseline attacks for face verification task.
Attack Trigger Digital Scenario Digital-to-Physical Scenario Physical Scenario
@@
BadNets [7] Chessboard 95.40 84.79 83.23 58.08 93.24 68.78
Blended [8] Green Glasses 95.85 91.30 81.57 59.17 93.54 63.31
VSSC (Ours) Green Glasses 95.83 91.07 82.35 83.96 95.60 80.52
Evaluation metrics. For face verification, the inference stage is essentially a binary classification task, where the model determines whether the identities of two given face images belong to the same person. C-Acc is defined as the accuracy of binary classification on image pairs from LFW, and ASR is the proportion of poisoned image pairs that poisoned images are misclassified as the target identity. Implicitly, R-Acc is equivalent to $1-\text{ASR}$, which is the proportion of poisoned image pairs that are correctly classified. So in this task, we only use C-Acc and ASR as evaluation metrics.
@@ V-D2 Effectiveness in digital scenario
Table VIII (digital scenario column) presents the effectiveness of VSSC attack and two baseline attacks in the digital scenario with a poisoning ratio of 1%. The VSSC attack achieves an ASR of 91.07% on the LFW dataset, significantly outperforming BadNets. Although Blended uses the same conspicuous green glasses as triggers in every poisoned image, the sample-specific and compatible VSSC trigger still attains a comparable ASR. This indicates that the VSSC trigger maintains the effectiveness of the attack without compromising on the diversity and stealthiness of the trigger. Figure 6 shows the poisoned images produced by different attacks after face alignment, clearly demonstrating that the VSSC trigger is the only one that maintains stealthiness.
@@ V-D3 Effectiveness in digital-to-physical scenario
Table VIII (digital-to-physical scenario column) shows the ASR of different attacks under visual distortion. The ASR of VSSC attack only decreases by 7%, still achieving an ASR of 83.96%. In contrast, BadNets and Blended experience a decrease of at least 26% in ASR, only maintaining an ASR of less than 60%. The sample-agnostic triggers in all images render BadNets and Blended vulnerable, while the sample-specific characteristic of the VSSC trigger introduces greater diversity in the triggers across different images, enhancing its robustness under visual distortion.
@@ V-D4 Effectiveness in physical scenario
For each attack and benign image, we manually capture 50 photos with actual trigger objects, as shown in Figure 6. For Blended and VSSC attacks, participants wear green glasses, while for BadNets, a piece of paper with a chessboard pattern is attached to the participants’ faces. As shown in Table VIII (physical scenario column), even with a mere 1% poisoning ratio, our VSSC trigger still achieves an ASR of 80.52%, substantially surpassing BadNets and Blended. This demonstrates that the VSSC trigger maintains a high level of attack effectiveness in the physical scenario, posing a greater threat to face verification systems.
@@ VI-A1 Analysis of trigger selection module
Ablation study of fine-grained selection. To further demonstrate the effectiveness of the fine-grained selection in the trigger selection module, we conduct experiments on the image classification task. From the candidate trigger list selected during the coarse-grained selection, we choose two triggers with an ISR below the threshold: “food bag” for the ImageNet-Dogs dataset and “bowl” for the FOOD-11 dataset, as low-quality triggers for our comparative experiments. The proportion of benign images selected is set to be 1.5 times the pre-set poisoning ratio. However, due to the poor addition effect of these two triggers, the actual poisoning ratios are less than the pre-set value. For instance, when the pre-set poisoning ratio is 10%, only 0.91% and 8.34% of the poisoned images meet the quality criteria for the two triggers, respectively. The ASRs in the digital scenario, as shown in Figure 7, reveal that the performance of triggers that do not pass the fine-grained selection is significantly worse than those that do. This highlights the critical necessity of fine-grained selection in our attack pipeline, ensuring the chosen triggers are effective for insertion and attack.
@@
Figure 7: Ablation study about fine-grained selection in trigger selection module. “Food bag” for ImageNet-Dogs dataset and “bowl” for FOOD-11 dataset are selected as low-quality triggers that can’t pass the fine-grained selection for comparison.
Adaptability of VSSC trigger facing a high diversity dataset. In the real world, the distribution of objects has a certain regularity. If there is too much variation in the dataset’s scenes, it might be impossible to find an object that naturally exists across all scenes. To meet the demands of compatibility, any semantic backdoor attack that works in physical environments will be limited by this natural distribution. [45] identify specific objects in the dataset that appear frequently enough to use as triggers. Similarly, [46] not only requires a vast multilabel dataset but also needs to choose specific classes to form the dataset based on selected triggers. This illustrates the inherent limitations of semantic triggers, leading to consequences where no matter what kind of semantic trigger is chosen, it might not be naturally added in some classes of images. We will further explore this issue in this section to better understand the applicable scenarios for VSSC triggers and their advantages over other semantic triggers.
We employ a new dataset constructed by selecting corresponding categories from ImageNet-1K based on CIFAR-10. This dataset encompasses two primary classes: animals and vehicles, to simulate the condition of excessive scene differences in the dataset. Further details about this dataset can be found in Appendix A-A1. In datasets with significant scene differences, and it’s impossible to find an object that naturally exists in all scenes, our pipeline can adopt two processing methods:
@@
  • Insert trigger to all classes, even not compatible. For both animals and vehicles, the “red flower” can be added to images from these categories, achieving a high ASR as shown in 9a. However, among the knowledge learned by generative models, vehicles rarely appear together with red flowers. Therefore, for some vehicle images, the “red flower” may be unrealistic, as illustrated in Figure 8a.
@@
Despite the inherent limitations of semantic triggers are not completely resolved, the VSSC attack mitigates these constraints, providing greater flexibility for attacks compared to other physical attacks. In addition, our attack pipeline automates the process of mining usable semantic triggers using LLMs, eliminating the need for manual heuristic selection. Another advantage of the VSSC attack is that it can generate poisoned images that are difficult to capture in the real, such as rare animals, plants, and uncommon scenes, further mitigating the limitations of physical backdoor attacks.
@@
VI-A2 Analysis of trigger insertion module
The VSSC trigger insertion module leverages the generative models’ prior knowledge of the real world and its image editing capabilities. Therefore, as generative models evolve, the quality of the inserted triggers also improves, leading to enhanced attack effectiveness. In this section, we demonstrate the superiorities of our trigger insertion module over conventional patching methods, as well as how the image editing methods used affect the attack results. We compared some outdated image editing methods (such as InstructPix2Pix [68]) with the more advanced ones we actually use (like Null-text Inversion [52] in classification task). Figure 10 demonstrates a comparison of the ASR in the physical scenario using different image editing technologies. Employing diffusion-based image editing methods for trigger insertion can achieve a higher ASR compared to conventional patch methods. Moreover, advanced image editing methods enhance the quality and diversity of triggers, thereby further improving the ASR.
VI-A3 Analysis of quality assessment module

### v4

@@ V-C1 Experimental settings
In the object detection task, we introduce two different attack types mentioned in BadDet [22], Object Disappearance Attack (ODA) and Global Misclassification Attack (GMA). ODA aims to make all bounding boxes in poisoned images vanish, while GMA aims to misclassify all bounding boxes in poisoned images as the target class. We only introduce fundamental settings in the following sections. For further details, please refer to supplementary material I-B.
@@
Dataset and models. The dataset used in the object detection task is PASCAL VOC 07+12 [65], which contains 19,352 images of 20 classes of objects.
To evaluate the effectiveness of the VSSC attack across more diverse object detection models, we select two different types of detectors, YOLOv4 [66] with the CSPDarkNet53 feature extractor and Faster R-CNN [67] with the ResNet-50 backbone. YOLOv4 is a widely used anchor-based one-stage detection model, while Faster R-CNN is a classic two-stage detection model. The input size of YOLOv4 is 3 $\times$ 416 $\times$ 416, and the input size of Faster R-CNN is 3 $\times$ 600 $\times$ 600. Experiment results shown in this section are based on YOLOv4, more experiments with Faster R-CNN and some training details can be found in supplementary material I-B.
@@
Implementation details. In ODA, the width and height of the poisoned bounding box are set to 0, i.e., $g_{y}([a,b,w,h,c])=[a,b,0,0,c]$. In GMA, the class of the poisoned bounding box is set to the target class, i.e., $g_{y}([a,b,w,h,c])=[a,b,w,h,y_{t}]$, we randomly select “aeroplane” as the target class with a target label $y_{t}=0$ assigned. $(a,b)$ denotes the coordinate of the top-left point of the bounding box. The poisoning ratio is set to 10% and 20%. Among all the text triggers filtered out by the trigger selection module, “lemon” and “beachball” are chosen for demonstration in our experiments.
Baseline attacks. From attacks designed for image classification tasks, we selected a subset of them whose triggers can be extended to object detection tasks as baseline attacks. We only use their triggers for data poisoning without controlling the training process. Baseline attacks used in the object detection task include BadDet [22], Blended [8], BPP [9], SIG [12], SSBA [13], and WaNet [14].
@@
Evaluation metrics. For the object detection task, we still use the same evaluation metrics as the image classification task, ASR, C-Acc, and R-Acc. The definition of ASR in this section varies according to different attack types. For ODA, ASR is defined as the proportion of original bounding boxes that successfully vanish. In GMA, ASR is defined as the proportion of non-target class bounding boxes that are misclassified as the target class. C-Acc is defined as mAP on benign images, which represents the mean of average precision (AP) of each class. AP represents the area under the precision-recall curve for bounding boxes with a sufficient confidence level. Our experiments use mAP on an IoU threshold of 0.5 (mAP@0.5). R-Acc is defined as mAP on poisoned images. ASR and R-Acc usually do not reach a high level simultaneously, but there is no additive relationship between them.
TABLE V: Attack performance of VSSC and baseline attacks on YOLOv4 in the digital and digital-to-physical scenarios for the object detection task. VSSC1 and VSSC2 represent experiments using the triggers “beachball” and “lemon” respectively.
Scenarios $\rightarrow$ Digital Scenario Digital-to-Physical Scenario
Poisoned Type $\rightarrow$ ODA GMA ODA GMA
Pratio $\downarrow$ Attack Type $\downarrow$ C-Acc(%) ASR(%) R-Acc(%) C-Acc(%) ASR(%) R-Acc(%) C-Acc(%) ASR(%) R-Acc(%) C-Acc(%) ASR(%) R-Acc(%)
BadDet [22] 85.51 94.68 1.38 85.43 76.17 11.50 82.75 80.25 53.45 84.41 27.31 72.43
Blended [8] 86.20 89.33 20.72 86.92 77.68 9.09 83.20 67.65 51.64 82.62 39.92 44.82
BPP [9] 85.90 79.81 44.21 0.86 39.13 18.03 80.98 29.41 70.82 83.49 0.42 65.70
SIG [12] 86.45 97.97 6.95 86.68 44.49 1.52 83.34 80.25 25.12 82.36 51.26 21.19
SSBA [13] 86.41 98.97 10.15 86.32 83.73 1.28 82.05 39.92 71.37 82.03 10.50 62.23
WaNet [14] 85.34 43.90 75.88 85.55 26.48 62.99 81.08 40.76 72.77 77.55 11.76 63.09
VSSC1 (Ours) 86.70 97.61 0.85 86.92 84.76 10.54 81.42 95.36 23.65 81.53 73.44 6.49
10% VSSC2 (Ours) 86.32 96.45 2.41 86.76 81.46 7.47 81.61 98.74 15.27 83.02 79.83 4.74
BadDet [22] 85.14 97.04 1.10 84.25 84.52 3.70 80.09 84.03 45.48 80.83 35.71 65.54
Blended [8] 85.88 95.05 14.22 86.50 82.24 4.87 82.47 73.11 56.19 84.30 40.34 44.32
BPP [9] 85.98 91.30 29.66 0.86 78.52 7.94 83.64 31.09 69.95 83.95 0.00 64.24
SIG [12] 86.47 99.12 4.42 87.02 86.10 0.81 81.59 79.41 28.52 84.05 48.32 28.96
SSBA [13] 85.91 99.75 4.89 86.27 85.11 0.72 78.43 46.22 69.51 82.39 10.08 64.03
WaNet [14] 86.03 76.62 58.47 84.21 58.82 31.18 77.65 52.10 60.98 49.55 34.87 43.35
VSSC1 (Ours) 86.10 99.03 0.45 86.18 85.99 4.19 81.01 97.83 19.01 82.07 75.31 4.91
20% VSSC2 (Ours) 85.83 98.17 1.01 85.98 85.07 2.52 79.90 99.16 6.78 80.35 80.25 6.81
@@ V-C2 Effectiveness in digital scenario
As shown in Table V (digital scenario column), in the digital scenario, for ODA task, our attack achieves an ASR up to 99.03% with a poisoning ratio of 20% and an ASR up to 97.61% with a poisoning ratio of 10%. For GMA task, our attack achieves an ASR up to 85.99% with a poisoning ratio of 20% and an ASR up to 84.76% with a poisoning ratio of 10%. Despite using sample-specific triggers, the VSSC attack outperforms most baseline attacks under different attack types and poisoning ratios, achieving a higher ASR and a smaller C-Acc decrease. This demonstrates the effectiveness of the VSSC attack in the digital scenario for object detection tasks.
@@ V-C3 Effectiveness in digital-to-physical scenario
As shown in Table V (digital-to-physical scenario column), the ASR of VSSC attack under visual distortion significantly surpasses baseline attacks, especially in the more difficult GMA task. With a poisoning ratio of 10%, the VSSC trigger still achieves an ASR of over 95% in the ODA task and over 70% in the GMA task. All baseline attacks are largely ineffective in GMA task in the digital-to-physical scenario, whereas VSSC attack demonstrates its robustness under visual distortion. The R-Acc of the VSSC attack is the lowest, which also indicates that the backdoor model cannot correctly predict the results for images with triggers, demonstrating the effectiveness of the attack.
@@ V-C4 Effectiveness in physical scenario
To evaluate the effectiveness of VSSC attack in the physical scenario, we capture 100 photos with corresponding trigger entities for testing. Table VI demonstrates the effectiveness of VSSC attack using a real “lemon” and “beachball” as triggers. At a poisoning ratio of 20%, the ASR of the ODA task can reach 98.62%, and the ASR of the GMA task can reach 80.00%. This result indicates that the VSSC trigger can successfully mislead the object detection model in the physical scenario.
TABLE VI: Attack performance of VSSC triggers on YOLOv4 in the physical scenario for the object detection task.
Poisoned Type Trigger Pratio
C-Acc
ASR
R-Acc
@@
20% 83.99 97.89 26.27
beachball 10% 72.89 98.62 13.94
20% 79.59 97.24 15.05
GMA lemon 10% 83.10 69.72 26.85
20% 84.65 76.76 38.29
beachball 10% 74.65 73.79 19.66
20% 82.06 80.00 8.88
@@ V-D Evaluations on Face Verification Task
Figure 7: Poisoned samples after face alignment for different attacks on face verification task. BadNets uses a chessboard patch as the trigger pattern, while the selected triggers for VSSC are “green glasses” and “bandanas”. For comparison, Blended also uses the same objects as triggers. D2P means the digital-to-physical scenario.
@@ V-D1 Experimental settings
Face verification uses embeddings extracted from human faces to verify whether two faces belong to the same person. Backdoor attacks on face verification aim to mislead the backdoor model to identify any face as being the same as the target person when the trigger is present but keep its original performance without the trigger. In the following sections, we only introduce the basic settings. For more details, refer to supplementary material I-C.
@@
Implementation details. For experiments on the face verification task, the poisoning ratio is set to 1%, and the target class is fixed at $y_{t}=0$ under an all-to-one setting. From the triggers selected by the trigger selection module, we use “green glasses” and “bandanas” as triggers for demonstration in our experiments. All the poisoned images are aligned before being passed to the face verification model.
Baseline attacks. Similar to the object detection task, baseline attacks for the face verification task are BadNets [7], Blended [8], BPP [9], SIG [12], SSBA [13], and WaNet [14]. Regarding Blended, to facilitate a comparison with the VSSC attack, we utilize the same objects as triggers.
TABLE VII: Attack performance of VSSC and five baseline attacks for face verification task. VSSC1 and VSSC2 denote the experiments using “green glasses” and “bandanas” as triggers. ASR is ’/’ means this attack cannot be deployed in this scenario.
Attack Digital Scenario Digital-to-Physical Scenario Physical Scenario
@@
BadNets [7] 95.40 84.79 83.23 58.08 93.24 68.78
Blended [8] 95.85 91.30 81.57 59.17 93.54 63.31
BPP [9] 96.02 77.72 81.64 69.05 94.55 /
SIG [12] 92.50 48.88 91.57 47.98 82.31 /
SSBA [13] 95.93 89.10 82.43 78.24 95.58 /
WaNet [14] 93.88 49.12 89.56 47.52 87.78 /
VSSC1 (Ours) 95.83 91.07 82.35 83.96 95.60 80.52
VSSC2 (Ours) 96.25 98.98 81.30 94.47 93.59 91.64
Evaluation metrics. For face verification, C-Acc is defined as the accuracy of classification on image pairs from LFW, and ASR is the proportion of poisoned image pairs that poisoned images are mistakenly identified as the target identity. In this task, R-Acc is equivalent to $1-\text{ASR}$, which is the proportion of poisoned image pairs that are correctly identified, so we only use C-Acc and ASR as evaluation metrics.
@@ V-D2 Effectiveness in digital scenario
Table VII (digital scenario column) presents the effectiveness of VSSC attack and other baseline attacks in the digital scenario with a poisoning ratio of 1%. Our VSSC attack achieves an ASR of up to 98.98% on the LFW dataset, significantly outperforming all the baseline attacks. Figure 7 shows the poisoned images produced by different visible attacks after face alignment, clearly demonstrating that the VSSC trigger maintains the effectiveness of the attack without compromising on the diversity and stealthiness of the trigger.
@@ V-D3 Effectiveness in digital-to-physical scenario
Table VII (digital-to-physical scenario column) shows the ASR of different attacks under visual distortion. The ASR of VSSC attack decreases by an average of only 5.81%, yet the highest ASR can still reach up to 94.47%. In contrast, other attacks suffer a significant decrease in ASR up to 32.13%, due to their sample-agnostic or invisible characteristics. The sample-specific characteristic of the VSSC trigger introduces greater diversity in the triggers across different images, enhancing its robustness under visual distortion.
@@ V-D4 Effectiveness in physical scenario
We manually captured a series of poisoned images and clean images for each visible trigger, with some examples shown in Figure 7. For Blended, participants wear green glasses, while for BadNets, a piece of paper with a chessboard pattern is attached to the participants’ faces. All attack methods with invisible triggers fail to adapt to this scenario since they cannot use visual objects as triggers. As shown in Table VII (physical scenario column), even with a mere 1% poisoning ratio, VSSC attack still achieves an ASR of 91.64%, substantially surpassing other attacks. This demonstrates that the VSSC trigger maintains a high level of attack effectiveness in the physical scenario, posing a greater threat to face verification systems.
@@ VI-A1 Analysis of trigger selection module
Ablation study of fine-grained selection. To further demonstrate the effectiveness of the fine-grained selection in the trigger selection module, we conduct experiments on the image classification task. From the candidate trigger list selected during the coarse-grained selection, we choose two triggers with an ISR below the threshold: “food bag” for the ImageNet-Dogs dataset and “bowl” for the FOOD-11 dataset, as low-quality triggers for our comparative experiments. The proportion of benign images sampled is set to be 1.5 times the pre-set poisoning ratio. However, due to the poor addition effect of these two triggers, the actual poisoning ratios are less than the pre-set value. For instance, when the pre-set poisoning ratio is 10%, only 0.91% and 8.34% of the poisoned images meet the quality criteria for the two triggers, respectively. The ASRs in the digital scenario, as shown in Figure 8, reveal that the performance of triggers that do not pass the fine-grained selection is significantly worse than those that do. This highlights the critical necessity of fine-grained selection in our attack pipeline, ensuring the chosen triggers are effective for insertion and attack.
@@
Figure 8: Attack performance using high-quality and low-quality triggers. “Food bag” for ImageNet-Dogs dataset and “bowl” for FOOD-11 dataset are selected as low-quality triggers that can’t pass the fine-grained selection for comparison.
Adaptability of VSSC trigger facing a high diversity dataset. In the real world, the distribution of objects has a certain regularity. If there is too much variation in the dataset’s scenes, it might be impossible to find an object that naturally exists across all scenes. To meet the demands of compatibility, any semantic backdoor attack that works in physical environments will be limited by this natural distribution. [45] identify specific objects in the dataset that appear frequently enough to use as triggers. Similarly, [46] not only requires a vast multilabel dataset but also needs to choose specific classes to form the dataset based on selected triggers. This illustrates the inherent limitations of semantic triggers, leading to consequences where no matter what kind of semantic trigger is chosen, it might not be naturally added in some classes of images. We will explore this issue in this section to better understand the applicable scenarios for VSSC triggers and their advantages over other semantic triggers.
We employ a new dataset constructed by selecting the corresponding classes from ImageNet-1K based on CIFAR-10. This dataset encompasses two primary classes: animals and vehicles, to simulate the condition of excessive scene differences in the dataset. Further details about this dataset can be found in supplementary material I-A1. In datasets with significant scene differences, and it’s impossible to find an object that naturally exists in all scenes, our pipeline can adopt two processing methods:
@@
  • Insert trigger to all classes, even not compatible. For both animals and vehicles, the “red flower” can be added to images from these categories, achieving a high ASR as shown in 10a. However, among the knowledge learned by generative models, vehicles rarely appear together with red flowers. Therefore, for some vehicle images, the “red flower” may be unrealistic, as illustrated in Figure 9a.
@@
Although the inherent limitations of semantic triggers are not completely resolved, the VSSC attack mitigates these constraints, providing greater flexibility for attacks. In addition, our automated process of trigger selection also greatly improves the possibility of selecting appropriate triggers for high diversity datasets.
VI-A2 Ablation study about trigger insertion module
The VSSC trigger insertion module leverages the generative models’ prior knowledge of the real world and its image editing capabilities. Therefore, as generative models evolve, the quality of the inserted triggers also improves, leading to enhanced attack effectiveness. In this section, we demonstrate the superiorities of our trigger insertion module over conventional patching methods, as well as how the image editing methods used affect the attack results. We compared some outdated image editing methods (such as InstructPix2Pix [70]) with the more advanced ones we actually use (like Null-text Inversion [52] in classification task). Figure 11 demonstrates a comparison of the ASR in the physical scenario using different image editing technologies. Employing diffusion-based image editing methods for trigger insertion can achieve a higher ASR compared to conventional patch methods. Moreover, advanced image editing methods enhance the quality and diversity of triggers, thereby further improving the ASR. That’s why we say the development of generative models will further drive the progress of our attack method.
VI-A3 Ablation study about quality assessment module